OpenStudy (anonymous):

find the area bounded by the y axis and the curves y=e^2x and y =e^X+2

OpenStudy (anonymous):

well we already know the region i guess

OpenStudy (anonymous):

is it \[\int\limits_{1}^{3} \log2 .dy\] ?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

you have to integrate from the y - axis which is x = 0 to the point of intersection of the curves

OpenStudy (anonymous):

which was the previous exercise

OpenStudy (anonymous):

you already found where these intersect. that is your upper limit of integration

OpenStudy (anonymous):

lower limit is 0 since you are starting at the y - axis

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

take bigger function minus smaller one on this region

OpenStudy (anonymous):

\[\int_0^{\ln(2)} e^x+2-e^{2x}dx\] is the integral you want

OpenStudy (anonymous):

you mean e^x -2-e^2x?

OpenStudy (anonymous):

no you wrote \[y=e^{2x}\] and \[y=e^x+2\] yes? it is bigger one minus smaller one. on the interval \[(0,\ln(2))\] \[e^x+2\] is larger, so it is just what i wrote \[\int e^x+2-e^{2x}dx\]

OpenStudy (anonymous):

oh i understand, my apologies

OpenStudy (anonymous):

now find the anti derivative, plug in ln(2) plug in 0 and subtract. done

OpenStudy (anonymous):

what answer did you get cause i got 2 and according to the answers its wrong

OpenStudy (anonymous):

oh wait actuall nvm

OpenStudy (anonymous):

disregard what i just said

OpenStudy (anonymous):

antiderivative of \[e^x\] is \[e^x\] antiderivative of \[2\] is \[2x\] antiderivative of \[-e^{-2x}\] is \[\frac{1}{2}e^{-2x}\]

OpenStudy (anonymous):

you have \[e^x+2x+\frac{1}{2}e^{-2x}|_0^{\ln(2)}\]

OpenStudy (anonymous):

isn't it 1/2e^2x

OpenStudy (anonymous):

i mean sorry -e^2x

OpenStudy (anonymous):

yes i don't know where i got the -x from

OpenStudy (anonymous):

it is \[-\frac{1}{2}e^{2x}\]

OpenStudy (anonymous):

so it should be \[e^x+2x-\frac{1}{2}e^{2x}|_0^{\ln(2)}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i get \[e^{\ln(2)}+2\ln(2)-\frac{1}{2}e^{\ln(2)}-(e^0-\frac{1}{2}e^0)\]

OpenStudy (anonymous):

which cleans up to \[2+2\ln(2)-1-1+\frac{1}{2}\] \[=2\ln(2)+\frac{1}{2}\] if my arithmetic is correct

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

made some mistake somewhere

OpenStudy (anonymous):

oh i see it. i should be \[e^\ln(2)+2\ln(2)-\frac{1}{2}e^{2\times \ln(2)}-(e^0-\frac{1}{2}e^0)\]

OpenStudy (anonymous):

now it is right. get \[2+2\ln(2)-2-1+\frac{1}{2}\] \[=2\ln(2)-\frac{1}{2}\]

OpenStudy (anonymous):

YES YOUR ANSWER IS CORRECT ACCORDING TO THE SOLUTION PAPER !

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

that is the correct one. i was off by 1 because if forgot it was \[e^{2x}\]

OpenStudy (anonymous):

yw